OpenStudy (anonymous):

Evaluate the integral \[\int\limits_{}^{}x/\sqrt{1-x^4}\]

OpenStudy (anonymous):

\[\frac{\text{ArcSin}\left[x^2\right]}{2}+c \]

myininaya (myininaya):

triangle problem!

myininaya (myininaya):

well there are formulas you can use but i like using a right triangle for this

myininaya (myininaya):

let cos(theta)=x^2

OpenStudy (anonymous):

gotta use trig sub i think but i get cot

OpenStudy (zarkon):

did you try myinimaya's substitution? or \[\sin(\theta)=x^2\] both will work

myininaya (myininaya):

yes both will like totally work :)

myininaya (myininaya):

if you giveup look at my attachment

OpenStudy (anonymous):

wow u make it look so easy

OpenStudy (zarkon):

the smart ones do that

myininaya (myininaya):

i like not remembering those formulas for this in fact i don't remember them so i do a substitution for these all the time

OpenStudy (anonymous):

there is another nice one for you

OpenStudy (anonymous):

@cinnamon, you know the best way to do them (besides computer algebra system like wolfram)?

OpenStudy (anonymous):

look in the back of the text they have all the formulas there

OpenStudy (anonymous):

back cover. i bet i can find this one

OpenStudy (anonymous):

ok i am still looking but i bet it is here

OpenStudy (anonymous):

ok i can't find it. but before you do a trig sub try a u-sub \[u=x^2\] \[du=2xdx\] to get \[\frac{1}{2}\int \frac{du}{\sqrt{1-u^2}}\]

OpenStudy (anonymous):

i gotta write this stuff out for my final

OpenStudy (anonymous):

this gets \[\frac{1}{2}\sin^{-1}(u)\] right away

OpenStudy (anonymous):

so answer is \[\frac{1}{2}\sin^{-1}(x^2)\]